write an equation in point slope form for the line with a slope of -9 that passes through (1,-3)
Answer by ikleyn(53751)

.

In the most general case, the line, which has the slope "m" and goes through the point (a,b), has an equation


    y - b = m*(x-a).


In your case,  m= -9,  a= 1,  b= -3, so the line has an equation


    y - (-3) = (-9)*(x-1),   or

    y + 3 = -9x + 9,   or

    y = -9x + 6


or any other equivalent equation.

Solved, answered and explained.

---------------

As a bonus for you,  consider the lessons in this site
    - Find the slope of a straight line in a coordinate plane passing through two given points
    - Equation for a straight line having a given slope and passing through a given point
    - Solving problems related to the slope of a straight line
    - Equation for a straight line in a coordinate plane passing through two given points
    - Equation for a straight line parallel to a given line and passing through a given point
    - Equation for a straight line perpendicular to a given line and passing through a given point
that are closely related to your problem.

You have a happy and rare opportunity to learn the subject from these lessons, written in compact and clear form.

Consider them as your textbook,  handbook,  guidebook,  tutorials and  (free of charge)  home teacher.